# Quantum optical coherence theory based on Feynman’s path integral

### v1

@@ II.2.2 Distinguishable and indistinguishable paths
Indistinguishable paths is a unique concept in quantum theory. All the paths in classical theory are distinguishable. One reason is that measurement are different in quantum and classical theories Wheeler and Zurek (2014). In classical theory, measurement precision can be arbitrarily high (at least in principle) without introducing unpredictable influence to the measured system. In quantum theory, measurement precision can not be arbitrarily high due to the measurement will introduce unpredictable influence on the measured system. Another reason is that there is no identical particles in classical theory. As Griffiths and Schroeter mentioned in their book that all electrons are utterly identical, in a way that no two classical objects can never be, identical particles is a unique concept in quantum theory Griffiths and Schroeter (2018).
Figure 2: Distinguishable and indistinguishable paths. S1 and S2 are two sources can emit quantum or classical particles. D is a single-particle detector.
Taking Fig. 2 as an example to illustrate distinguishable and indistinguishable paths. S1 and S2 are two sources that can emit either quantum or classical particles. D is a single-particle detector. If the detector D detects a particle. The particle comes from either S1 or S2. In classical theory, the particles emitted by these two sources are always distinguishable. Even if all the properties such as mass, volume, color and so on, are identical [^1] This assumption is impossible in classical theory. For instance, two particles having the same weight in classical theory is only valid in some precision. Since classical particles are infinitely divisible, one can always find higher precision in which these two particles having different weights., the momentum difference between these two cases can be employed to distinguish these two paths. Different paths are always distinguishable in classical theory. Things become different in quantum theory. Assuming the particles emitted by S1 and S2 are identical. The only difference between these two paths is the momentum difference. The measurement precision of momentum in quantum theory is limited by Heisenberg uncertainty principle Heisenberg (1949)
@@
where $\Delta p_{x}$ and $\Delta x$ are the uncertainties of momentum and position of the measured particle, respectively. The uncertainty of position is usually determined by the detection mechanism of the employed detector. For instance, the position uncertainty of photon detection is usually at the size of atom absorbing the photons via photoelectric effect, which is usually at $10^{-10}$ m range for metal atoms Shackelford (2016). The paths with momentum difference less than $h/10^{-10}$ are indistinguishable. Different paths can be indistinguishable in quantum theory. Heisenberg uncertainty principle should be employed to analyze the path difference when detail measuring apparatus is taken into account.
II.3 What is photon?
The concept of photon can trace back to Planck Plank (1901) and Einstein Einstein (1905). The answer to the question of what is photon is still not satisfactory. However, it does not stop photons appear everywhere even when there is no need for them Schleich (2003). As stated by Einstein in 1951: All these fifty years of conscious brooding have brought me no nearer to the answer to the question, -What are light quanta? Nowadays every Tom, Dick and Harry thinks he knows it, but he is mistaken Afshar et al. (2007). Despite significant progress in our understanding, it remains an open question Roychoudhuri and Roy (2003). In quantum optical coherence theory based on Feynman’s path integral, the concept of photon is unavoidable. It is impossible for us to answer the question of what is photon right now. However, it is helpful to clarify what we are talking about when the concept photon is employed.
We think photon is a particle, which is described by photon wave function $\psi(\vec{r},t)$. The probability to find a photon at $(\vec{r},t)$ in a volume $dV$ is given by $|\psi(\vec{r},t)|^{2}dV$. The initial phase of a photon is determined by the emission time.
@@ II.3.1 Photon is a particle
Wave-particle duality is well-accepted for photon. However, when we talk about photon in this paper, we try to picture it as a particle like electron, proton and so on. The particle is described by a photon wave function $\psi(\vec{r},t)$, which is different from the concept of particle in classical theory.
The evidences for that photon is a particle are photoelectric effect Hertz (1887) and photon antibunching effect Paul (1982). Hertz observed the following phenomena when shining light on the surface of a metal:
@@
(III) The time delay $T$ between the light shining on the metal and the ejected photoelectron is very short.
It is experimentally proved by Lawrence and Beams that $T$ is shorter than 3 ns Lawrence and Beams (1928). Forrester et al. further proved that $T$ is much shorter than 0.1 ns Forrester et al. (1955). The first two phenomena can be interpreted by wave theory. However, the last phenomenon can only be interpreted by treating photon is a particle Muthukrishnan et al. (2017).
The second evidence for that photon is a particle is photon antibuching effect Paul (1982). The physics of photon antibunching is that a photon can never trigger two separated single-photon detectors simultaneously. The key to observe photon antibunching is to generate single-photon state. Kimble et al. Kimble et al. (1977) and Grangier et al. Grangier et al. (1986) observed photon antibunching effect with different ways to generate single-photon state. Classical electromagnetic wave theory can not interpret photon antibunching effect Paul (1982); Klyshko (1996).
There are also other evidence for photon is a particle. For instance, when a beam of light with low intensity is incident to a interferometer, the first-order interference pattern is accumulated point by point on the observation plane Dimitrova and Weis (2008).
@@ II.3.2 Photon’s wave function
Strictly speaking, there is no wave function since for photon there is no position operator for photon Bohm (2012). However, one can define photon’s position operator from the point view of detection. Photon is usually detected by photoelectric effect. The position resolution is limited by the size of atom. As long as the position accuracy of photon’s wave function is not arbitrary high, one can define and use photon’s wave function by analogy of the wave function of massive particle.
@@
where $K$ is a normalization constant, $r=|\vec{r}-\vec{r}_{0}|$ is the distance between the position of the atom $\vec{r}_{0}$ and the position of photon $\vec{r}$, $\eta$ is the angle of $\vec{r}$ comparing to the electric dipole of atom, $\theta(t-\frac{r}{c})$ is a step function, which means that photon can only be detected when photon propagates at position $\vec{r}$, $\omega$ is the frequency of photon, $\Gamma$ is the width of up-level energy of two-level system, $\exp{[-i(\omega+i\frac{\Gamma}{2})(t-\frac{r}{c})]}$ means that the probability to detect the photon is decayed by $e^{-\Gamma t}$.
Equation (10) only gives the temporal part of photon’s wave function. The spatial part of photon’s wave function can be treated the same as the spatial structure of classical electric field Scully and Zubairy (1997). For instance, the electric field emitted by a point source is spherical wave in classical theory. The spatial wave function of single photon emitted by a point source is also spherical.
@@ II.3.3 Photon’s phase
Photon’s phase is defined by analogy of the phase of electric field in classical electromagnetic theory Dirac (1930). Classical electromagnetic field can be expressed as Born and Wolf (2013)
@@
The uncertainties of phase ($\Delta\varphi$) and number of photons ($\Delta N$) satisfies the following uncertainty principle Pegg and Barnett (1997)
@@
For number state $|N\rangle$, the uncertainty of photon number is zero, which means that the phase of photon in number state is complete uncertain. It can be proved that the phase distribution function,
@@
of single-photon state equals $1/2\pi$, which means the phase of photon in single-photon state is equally distributed between 0 and $2\pi$ Pegg and Barnett (1997).
Feynman assumed that the initial phase of photon is determined by the emission time Feynman (2006). Based on the phase relationship between different photons, light can be categorized into two groups. The first group is that the initial phases of photons in a light beam are random. The photons in this type of light are usually emitted by spontaneous emission. Chaotic light, thermal light, number state are examples of this type of light. The second group is that the initial phases of photons in a light beam are identical. The photons in this type of light are usually emitted by stimulated emission. Single-mode laser, amplified spontaneous emission belongs to the second type.
@@ II.4 Coherence volume and degeneracy factor of light
The key to apply superposition principle in Feynman’s path integral to calculate optical coherence is distinguishing different paths for an event are distinguishable or not. Heisenberg’s uncertainty principle should be employed to judge whether different paths are distinguishable or not. Indistinguishable path is close related to indistinguishable photons. Photons within the same coherence volume are indistinguishable Martienssen and Spiller (1964). The uncertainties of coordinates ($\Delta x$, $\Delta y$, and $\Delta z$) and momentum ($\Delta p_{x}$, $\Delta p_{y}$, and $\Delta p_{z}$) of photons within the same coherence volume satisfies the following equation,
$\Delta x\Delta p_{x}=h$
$\Delta y\Delta p_{y}=h$ (15)
@@
where $h$ is Planck constant, $v$ is the frequency of radiation, $k_{B}$ is Boltzmann constant, $T$ is temperature of blackbody. The degeneracy factor of usual thermal light is much less than 1. For instance, the degeneracy factor of sun light at 532 nm equals 0.0073. On the other hand, the degeneracy factor of laser light can be much larger than 1. For instance, the degeneracy factor a 1-mW single-mode continuous-wave He-Ne laser light equals $3\times 10^{9}$ Mandel and Wolf (1995).
@@ III First-order interference of light
The first-order interference of light can be well interpreted in classical optical coherence theory based on Maxwell’s electromagnetic theory Born and Wolf (2013). Dirac interpreted the first-order interference of light in the language of photon and gave the well-known statement about the first-order interference of light: “Each photon then interferes only with itself. Interferences between two different photons never occurs Dirac (1930).” The contents in this section are mainly based on Dirac’s statement. For a comprehensive review about the first-order interference of light, please refer to Refs. Born and Wolf (2013); Hecht (2012); Hariharan (2003).
@@ III.1 First-order interference of one light beam
The first-order interference of one light beam means that light emitted by one source is split into two or more components and then combined together to observe interference. It corresponds to that each photon interferes only with itself in Dirac’s statement. Taking Mach-Zehnder interferometer shown in Fig. 3 as an example to show how to employ Feynman’s path integral to discuss the first-order interference of one light beam.
Figure 3: Mach-Zehnder interferometer. S is light source. BS1 and BS2 are two non-polarizing 1:1 beam splitter. M1 and M2 are two mirrors. D is a single-photon detector.
The event for the single-photon detector D detecting a photon in Fig. 3 has two different paths. The first one is the detected photon taking the upper path (transmitting through beam splitter BS1, reflected by mirror M1, and reflected by beam splitter BS2) and recorded as path 1 for simplicity. The probability amplitude for photon taking path 1 is written as $A_{1}(S,D)$. The second path is the photon taking the lower path (reflected by BS1 and M2, transmitting through BS2) and recorded as path 2. The probability amplitude for photon taking path 2 is written as $A_{2}(S,D)$. If these two different paths to trigger a photon detection event are indistinguishable, the probability function for the $j$th detected photon is
$P_{j}(t)$ $=$ $|A_{1}(S,D)+A_{2}(S,D)|^{2}$ (17)
$=$ $|e^{i\varphi_{j}}K_{1}(S,D)+e^{i\varphi_{j}}K_{2}(S,D)|^{2},$
where $j$ is a positive integer, $\varphi_{j}$ is the initial phase of the $j$th detected photon, $K_{1}(S,D)$ is photon’s Feynman propagator from the source to detector via path 1, $A_{1}(S,D)=e^{i\varphi_{j}}K_{1}(S,D)$, $\varphi_{j}$ is the initial phase of the $j$th detected photon. Similar definition holds for $K_{2}(S,D)$. The first-order interference of one light beam is due to photon interferes with itself. The initial phases are the same for two different probability amplitudes.
The first-order interference pattern is proportional to the sum of the probability functions of all the detected photon,
@@
where $N$ is the number of detected photon. For a large enough number of photons,
@@
where delimited-⟨$\langle$..$\rangle$ … ⟩ is ensemble average by taking all possible realizations Mandel and Wolf (1995). Substituting Eq. (17) into Eq. (19), the first-order interference pattern in Fig. 3 is proportional to
@@
The initial phase of photon, $\varphi_{j}$, does not contribute to the first-order interference pattern in Fig. 3. No matter whether thermal or laser light is employed, as long as the frequency bandwidth and spatial distribution are the same, the observed first-order interference pattern will be identical. The reason is that the first-order interference is a result of photon interferes with itself. The initial phases of different paths for the $j$th detected photon are identical.
For a point light source, photon’s Feynman propagator can be approximated as
@@
where $(\vec{r}_{0},t_{0})$ is the space-time coordinate of the photon being emitted, $(\vec{r},t)$ is the space-time coordinate of the photon being detected, $\omega_{0}$ is the angular frequency, $\vec{k}$ is wave vector. For simplicity, single-frequency is assumed in the following calculations. Substituting Eq. (21) into Eq. (20) and considering the temporal part, the first-order interference pattern in Fig. 3 is
@@
where $t_{1}$ and $t_{2}$ are the travailing time of photon from S to D via path 1 and 2, respectively. The first-order interference pattern is proportional to the probability function of the $j$th detected photon in Fig. 3, which is the same as the first-order interference pattern obtained in classical optical coherence theory Born and Wolf (2013).
@@
where $f(\omega)$ is the frequency distribution function, $\hat{a}^{\dagger}(\omega)$ is the creation operator, $|0\rangle$ is vacuum state.
When the time difference between the photon taking two paths, $|t_{1}-t_{2}|$, is larger than the coherence time of the employed single-photon state, the two different paths to trigger a photon detection event at D are distinguishable. The probability function for the $j$th detected photon equals
@@
$=$ $|e^{i\varphi_{j}}K_{1}(S,D)|^{2}+|e^{i\varphi_{j}}K_{2}(S,D)|^{2}$
@@
where $t$ is the time of photon being detected by detector, $t_{1}=t-t_{01}$, $t_{01}$ is the emitting time of photon detected by D via path 1, $t_{2}=t-t_{02}$ and $t_{02}$ is the emitting time of photon detected by D via path 2. The probability of the $j$th detected photon is the sum of the probabilities of the photon taking two different paths. There is no first-order interference.
When the time difference, $|t_{1}-t_{2}|$, is less than the coherence time of the employed single-photon state, these two different paths to trigger a photon detection event at D are indistinguishable. The probability function of the $j$th detected photon is
$P_{j}(t)$ $=$ $|e^{i\varphi_{j}}K_{1}(S,D)+e^{i\varphi_{j}}K_{2}(S,D)|^{2}$
@@
where $A$ is a normalization constant, $\sigma^{2}$ is variance of Gaussian function, $\omega_{0}$ is the central frequency of photon. Equation (III.1) can be simplified as Shih (2020)
@@
The first-order interference pattern can not be observed with only one photon. The first-order interference pattern of single-photon state can be observed by preparing many identical single-photon states and put these single photon states into the interferometer one by one. By collecting large enough number of photons, the first-order interference pattern with single-photon source in Fig. 3 is
@@
which is the same the probability of the $j$th detected photon.
Figure 4: First-order interference pattern of single-photon states in Mach-Zehnder interferometer
Figure 4 shows the first-order interference pattern of single-photon state with Gaussian spectrum in a Mach-Zehnder interferometer. $\tau=t_{1}-t_{2}$ is the time difference between path 1 and 2, I is the intensity of light. When the time difference equals 0, the visibility of the first-order interference pattern can reach 100%. As the time difference becomes larger, the visibility of first-order interference pattern decreases. When the time difference is larger than the coherence length of light, the visibility goes to zero, which is consistent with Eq. (III.1). The simulated result in Fig. 4 is consistent with experimental results with single photons Braig et al. (2003)
Similar method can be employed to calculate the spatial first-order interference pattern of one light beam in Feynman’s path integral.

### v2

@@ II.2.2 Distinguishable and indistinguishable paths
Indistinguishable paths are a unique concept in quantum theory. All the paths are distinguishable in classical theory. One reason is that measurements are different in quantum and classical theories Wheeler and Zurek (2014). In classical theory, measurement precision can be arbitrarily high (at least in principle) without introducing unpredictable influence to the measured system. In quantum theory, Measurement precision cannot be arbitrarily high because measurements will introduce unpredictable influence and the uncertainty of the measurement is limited by Heisenberg uncertainty principle. Another reason is that there are no identical particles in classical theory. As Griffiths and Schroeter mention in their book that all electrons are utterly identical, in a way that no two classical objects can never be, identical particle is a unique concept in quantum theory Griffiths and Schroeter (2018).
Figure 2: Distinguishable and indistinguishable paths. S1 and S2 are two sources which can emit quantum or classical particles. D is a single-particle detector.
Let us take the scheme shown in Fig. 2 as an example to illustrate distinguishable and indistinguishable paths. S1 and S2 are two sources that can emit either quantum or classical particles. D is a single-particle detector. If the detector D detects a particle. The particle comes from either S1 or S2. In classical theory, the particles emitted by these two sources are distinguishable. Even if all the properties such as mass, volume, color and so on, are identical [^1] This assumption is impossible in classical theory. For instance, two particles having the same weight in classical theory is only valid in some precision. Classical particles are infinitely divisible. One can always find higher precision in which these two particles having different weights., the momentum difference between these two cases can be employed to distinguish these two paths. Different paths are always distinguishable in classical theory. Things become different in quantum theory. Assuming the particles emitted by S1 and S2 are identical, the only difference between these two paths is the momentum difference. The measurement precision of momentum in quantum theory is limited by Heisenberg uncertainty principle Heisenberg (1949)
@@
where $\Delta p_{x}$ and $\Delta x$ are the uncertainties of momentum and position of the measured particle, respectively. The uncertainty of position is usually determined by the detection mechanism of the employed detector. Photon is usually detected by photoelectric effect. The position uncertainty of photon detection equals the size of the atom absorbing the photon, which is usually at $10^{-10}$ m range for metal atoms Shackelford (2016). The paths with momentum difference less than $h/10^{-10}$ are indistinguishable. Different paths can be indistinguishable in quantum theory. Heisenberg uncertainty principle is employed to analyze the path difference when detail measuring apparatus is taken into account.
II.3 What is a photon?
The concept of photon can trace back to Planck Plank (1901) and Einstein Einstein (1905). The answer to the question of what is a photon is still not satisfactory. However, it does not stop photons appear everywhere even when there is no need for them Schleich (2003). As stated by Einstein in 1951: All these fifty years of conscious brooding have brought me no nearer to the answer to the question, -What are light quanta? Nowadays every Tom, Dick and Harry thinks he knows it, but he is mistaken Afshar et al. (2007). Despite significant progress in our understanding, it remains an open question Roychoudhuri and Roy (2003). In quantum optical coherence theory based on Feynman’s path integral, the concept of photon is unavoidable. It is impossible for us to answer the question of what is a photon right now. However, it is helpful to clarify what we are talking about when the concept of photon is employed.
In short, we picture photon as a particle, which is described by photon wave function $\psi(\vec{r},t)$. The probability to find a photon at $(\vec{r},t)$ in a volume $dV$ is given by $|\psi(\vec{r},t)|^{2}dV$. The initial phase of a photon is determined by the emission time.
@@ II.3.1 Photon is a particle
Wave-particle duality is well-accepted for photon. When we talk about photon in this paper, we picture it as a particle like electron, proton and so on. The particle is described by a photon wave function $\psi(\vec{r},t)$, which is different from the concept of particle in classical theory.
The evidence that photon is a particle includes photoelectric effect Hertz (1887) and photon antibunching effect Paul (1982). Hertz observed the following phenomena when shining a light beam on the surface of a metal:
@@
(III) The time delay $T$ between the light shining on the metal and the ejected photo-electron is very short.
It is experimentally proved by Lawrence and Beams that $T$ is shorter than 3 ns Lawrence and Beams (1928). Forrester et al. further proved that $T$ is much shorter than 0.1 ns Forrester et al. (1955). The first two phenomena can be interpreted by wave theory. The last phenomenon can only be interpreted by treating photon is a particle Muthukrishnan et al. (2017).
The second evidence for that photon is particle is photon antibuching effect Paul (1982). The physics of photon antibunching is that one photon cannot trigger two separated single-photon detectors simultaneously. The key to observe photon antibunching is to generate single-photon state. Kimble et al. Kimble et al. (1977) and Grangier et al. Grangier et al. (1986) observed photon antibunching effect with different ways to generate single-photon state. Classical electromagnetic wave theory can not interpret photon antibunching effect Paul (1982); Klyshko (1996).
There are also other evidence for photon is a particle. For instance, when a light beam with low intensity is incident into an interferometer, the first-order interference pattern builds up point by point on the observation plane Dimitrova and Weis (2008).
@@ II.3.2 Photon’s wave function
Strictly speaking, there is no wave function for photon since there is no position operator for photon Bohm (2012). However, one can define photon’s position operator from the point of view of detection. Photon is usually detected by photoelectric effect. The position resolution is limited by the size of atom. As long as the position accuracy of photon’s wave function is lower than the size of atom, one can define and use photon’s wave function by analogy of the wave function of a massive particle.
@@
where $K$ is a normalization constant, $r=|\vec{r}-\vec{r}_{0}|$ is the distance between the position of the atom, $\vec{r}_{0}$, and the position of photon, $\vec{r}$, $\eta$ is the angle of $\vec{r}$ related to the electric dipole of atom, $\theta(t-\frac{r}{c})$ is a step function, which indicates that a photon can be detected at $\vec{r}$ only when it propagates to position $\vec{r}$, $\omega$ is the frequency of photon, $\Gamma$ is the width of the up-level energy of two-level system, $\exp{[-i(\omega+i\frac{\Gamma}{2})(t-\frac{r}{c})]}$ indicates that the probability of detecting the photon decays by $e^{-\Gamma t}$.
Equation (10) gives the temporal part of photon’s wave function. The spatial part of photon’s wave function can be treated the same as the spatial structure of classical electric field Scully and Zubairy (1997). For instance, the electric field emitted by a point source is spherical wave in classical theory. The spatial wave function of single photon emitted by a point source is also spherical.
@@ II.3.3 Photon’s phase
Photon’s phase is defined by analogy of the phase of electric field in classical electromagnetic theory Dirac (1930). Classical electric field can be expressed as Born and Wolf (2013)
@@
The uncertainties of phase ($\Delta\varphi$) and number of photons ($\Delta N$) satisfies the following uncertainty principle Pegg and Barnett (1997),
@@
For number state, $|N\rangle$, the uncertainty of photon number is zero, which means that the phases of photons in a number state are completely uncertain. It can be proved that the phase distribution function,
@@
of single-photon state equals $1/2\pi$, which means that the phase of the photon in a single-photon state is equally distributed between 0 and $2\pi$ Pegg and Barnett (1997).
Feynman assumed that the initial phase of a photon is determined by the emission time Feynman (2006). Based on the phase relationship between different photons in a light beam, light can be categorized into two groups. The first group is that the initial phases of photons in a light beam are random. The photons in this type of light are usually emitted by spontaneous emission. Chaotic light, thermal light, number state are examples of this type of light. The second group is that the initial phases of photons in a light beam are identical. The photons in this type of light are usually emitted by stimulated emission. Single-mode laser, amplified spontaneous emission belongs to the second type of light.
@@ II.4 Coherence volume and degeneracy factor of light
The key to applying superposition principle in Feynman’s path integral is to determine whether different paths for an event are distinguishable or not. Heisenberg uncertainty principle is employed to judge whether different paths are distinguishable or not. Indistinguishable paths are close related to indistinguishable particles. Photons within the same coherence volume are indistinguishable Martienssen and Spiller (1964). The uncertainties of coordinates ($\Delta x$, $\Delta y$, and $\Delta z$) and momentum ($\Delta p_{x}$, $\Delta p_{y}$, and $\Delta p_{z}$) of photons within the same coherence volume satisfies the following equations,
$\Delta x\Delta p_{x}=h,$
$\Delta y\Delta p_{y}=h,$ (15)
@@
where $h$ is the Planck constant, $\nu$ is the frequency of radiation, $k_{B}$ is the Boltzmann constant, $T$ is the temperature of blackbody. Thermal light can be treated as blackbody radiation. The degeneracy factor of usual thermal light is much less than 1. For instance, the degeneracy factor of sun light at 532 nm equals 0.0073. On the other hand, the degeneracy factor of laser light can be much larger than 1. For instance, the degeneracy factor a 1-mW single-mode continuous-wave He-Ne laser light beam equals $3\times 10^{9}$ Mandel and Wolf (1995).
@@ III First-order interference of light
The first-order interference of light can be well interpreted in classical optical coherence theory based on Maxwell’s electromagnetic theory Born and Wolf (2013). Dirac firstly interpreted the first-order interference of light in the language of photons and gave the well-known statement about the first-order interference of light: “Each photon then interferes only with itself. Interferences between two different photons never occurs Dirac (1930).” The contents in this section are mainly based on Dirac’s statement. For a comprehensive review about the first-order interference of light, please refer to Refs. Born and Wolf (2013); Hecht (2012); Hariharan (2003).
@@ III.1 First-order interference of one light beam
The first-order interference of one light beam means that light emitted by one source is split into two or more components and then combined together to observe the first-order interference pattern. It corresponds to each photon interferes only with itself in Dirac’s statement. Take Mach-Zehnder interferometer shown in Fig. 3 as an example to show how to employ Feynman’s path integral to discuss the first-order interference of one light beam.
Figure 3: Mach-Zehnder interferometer. S is a light source. BS1 and BS2 are two non-polarizing 1:1 beam splitter. M1 and M2 are two mirrors. D is a single-photon detector.
There are two different paths for a single-photon detector, D, detecting a photon in Fig. 3. The first path is the detected photon taking the upper path (transmitting through beam splitter BS1, reflected by mirror M1, and reflected by beam splitter BS2) and recorded as path 1 for simplicity. The probability for the photon taking the path is $1/2$ and the probability amplitude for the photon taking the path is $A_{1}(S,D)$. The second path is the photon taking the lower path (reflected by BS1 and M2, transmitting through BS2) and recorded as path 2. The probability for the photon taking the path is $1/2$ and the probability amplitude for photon taking the path is $A_{2}(S,D)$. If these two different paths to trigger a photon detection event are indistinguishable, the probability function for the $j$th detected photon is
$P_{j}(t)$ $=$ $|\frac{1}{\sqrt{2}}A_{1}(S,D)+\frac{1}{\sqrt{2}}A_{2}(S,D)|^{2}$ (17)
$\propto$ $|e^{i\varphi_{j}}K_{1}(S,D)+e^{i\varphi_{j}}K_{2}(S,D)|^{2},$
where $j$ is a positive integer, $\varphi_{j}$ is the initial phase of the $j$th detected photon, $K_{1}(S,D)$ is photon’s Feynman propagator from the source to the detector via path 1, $A_{1}(S,D)=e^{i\varphi_{j}}K_{1}(S,D)$. Similar definition holds for $K_{2}(S,D)$. The first-order interference of one light beam is due to photon interferes with itself. The initial phase is identical for the same photon taking two different paths.
The first-order interference pattern is proportional to the sum of the probability functions of all the detected photons,
@@
where $N$ is the number of the detected photons. For a large enough number of photons,
@@
where delimited-⟨$\langle$..$\rangle$ … ⟩ is ensemble average by taking all possible realizations Mandel and Wolf (1995); Shih (2020). Substituting Eq. (17) into Eq. (19), the first-order interference pattern in Fig. 3 is
@@
The initial phase of the photon, $\varphi_{j}$, does not contribute to the first-order interference pattern in Fig. 3. Regardless of whether thermal or laser light is used, as long as the frequency bandwidth and spatial distribution are identical, the observed first-order interference patterns will be the same. The reason is that the first-order interference is a result of photon interferes with itself.
For a point light source, the photon’s Feynman propagator can be approximated as
@@
where $(\vec{r}_{0},t_{0})$ is the space-time coordinate of the photon being emitted, $(\vec{r},t)$ is the space-time coordinate of the photon being detected, $\omega_{0}$ is angular frequency, $\vec{k}$ is wave vector. For simplicity, single-frequency light is assumed firstly. Substituting Eq. (21) into Eq. (20) and considering the temporal part, the first-order interference pattern in Fig. 3 is
@@
where $t_{1}$ and $t_{2}$ are the travailing time of the photon from S to D via path 1 and 2, respectively. The first-order interference pattern is proportional to the probability function of the $j$th detected photon in Fig. 3, which is the same as the first-order interference pattern obtained in classical optical coherence theory Born and Wolf (2013).
@@
where $f(\omega)$ is the frequency distribution function, $\hat{a}^{\dagger}(\omega)$ is the creation operator, $|0\rangle$ is vacuum state. When the time difference between the photon taking two paths, $|t_{1}-t_{2}|$, is longer than the coherence time of the employed single-photon state, the two different paths to trigger a photon detection event at D are distinguishable. The probability function for the $j$th detected photon equals
@@
$=$ $|\frac{1}{\sqrt{2}}e^{i\varphi_{j}}K_{1}(S,D)|^{2}+|\frac{1}{\sqrt{2}}e^{i\varphi_{j}}K_{2}(S,D)|^{2}$
@@
where $t$ is the time of photon being detected, $t_{1}=t-t_{01}$, $t_{01}$ is the emitting time of the photon detected by D via path 1, $t_{2}=t-t_{02}$ and $t_{02}$ is the emitting time of the photon detected by D via path 2. The probability function of the $j$th detected photon is the sum of the probability functions of the photon taking two different paths. There is no first-order interference.
When the time difference, $|t_{1}-t_{2}|$, is shorter than the coherence time of the employed single-photon state, these two different paths to trigger a photon detection event at D are indistinguishable. The probability function of the $j$th detected photon is
$P_{j}(t)$ $=$ $|\frac{1}{\sqrt{2}}e^{i\varphi_{j}}K_{1}(S,D)+\frac{1}{\sqrt{2}}e^{i\varphi_{j}}K_{2}(S,D)|^{2}$
@@
where $A$ is a normalization constant, $\sigma^{2}$ is variance of Gaussian function, $\omega_{0}$ is the central frequency. Equation (III.1) can be simplified as Shih (2020)
@@
The first-order interference pattern of light can not be observed with only one photon. The first-order interference pattern of a single-photon state can be observed by preparing many identical single-photon states and put these single-photon states into the interferometer one by one. By collecting large enough number of photons, the first-order interference pattern with single-photon states in Fig. 3 is
@@
which is the same the probability function of the $j$th detected photon.
Figure 4: First-order temporal interference pattern of single-photon states in a Mach-Zehnder interferometer. I is intensity of light and $\tau$ is time difference between photon taking these two paths.
Figure 4 shows the simulated first-order interference pattern of single-photon states with Gaussian spectrum in a Mach-Zehnder interferometer. $\tau=t_{1}-t_{2}$ is the time difference between paths 1 and 2, I is the intensity of light. When the time difference equals 0, the visibility of the first-order interference pattern can reach 100%. As the time difference becomes larger, the visibility of the first-order interference pattern decreases. When the time difference is larger than the coherence time of light, the visibility goes to zero, which is consistent with Eq. (III.1). The simulated result in Fig. 4 is consistent with the experimental results with single photons Braig et al. (2003)
Similar method can be employed to calculate the first-order spatial interference pattern of one light beam in Feynman’s path integral.